# Reframing Three-Dimensional Morphometrics Through Functional Data Innovations

### v1

@@ 1  Introduction
Functional data analysis (FDA) presents then a promising alternative by modeling curves as smooth functions within a functional space (Hilbert, Banach,… space) ramsay2005functional , treating the shape as a realisation of a stochastic process rather than a set of discrete coordinates. Srivastava and Klassen (2016)SrivastavaKlassen2016 motivate the need for rigorous statistical tools to analyse functions and shapes while handling invariances such as translation, rotation, scaling, and re-parameterisation. Their geometric framework unifies registration and shape analysis, enabling tools such as functional principal component analysis (FPCA) for vector-valued functions. Wu et al. (2024)wu2024shape presents a framework that integrates geometric shape analysis with functional data methods to better capture structural features of curves, such as peaks and valleys, beyond what traditional based approaches offer. The framework developed tools for shape fitting, shape-fPCA, and shape regression, extending these ideas to manifold-valued functions and demonstrating their effectiveness on real and simulated datasets.
@@
This work is organised as follows. Section 2 details the methods based on the GM and FDM frameworks. Section 3 presents a simulation study to compare FDM with its classical counterpart GM. Section 4 describes the application to 3D kangaroo landmark data. Section 6 concludes.
@@ 3.2  Simulation results
Figure 3 shows differences in classification performance, where functional data–driven approaches achieved higher predictive accuracy and better stability between folds. Figure 4 illustrates that functional pipelines reached the 95% variance threshold with fewer components than coordinate-based methods. The corresponding reconstruction plot (Figure Z) confirmed that the reconstructed shapes using the retained components preserved the curved structural information, thus validating the use of the 95% cutoff. Both elastic and soft-aligned pipelines outperformed the purely coordinate-based and standard functional approaches. The elastic variants generally achieved good accuracy and efficient variance capture, but their performance showed greater variability across folds, indicating reduced stability. In contrast, soft-aligned pipelines offered consistently better accuracy with reduced variability and clearer group separation in PC space (Figure 7).
@@
GM 9
FDM 3
Arc-GM 13
Arc-FDM 5
Elastic-SRV-FDM 4
Soft-SRV-FDM 3
Arc-Elastic-SRV-FDM 15
Arc-Soft-SRV-FDM 4
@@
GM
0.01994
FDM
0.00016
Arc-GM
0.02089
Arc-FDM
0.00425
Elastic-SRV-FDM
0.00261
Soft-SRV-FDM
0.00204
Arc-Elastic-SRV-FDM
0.00197
Arc-Soft-SRV-FDM
0.00192
@@
Figure 6: Figure 6: Reconstruction landmark configuration of specimen 1 via arc-elastic-SRV-FDM (original, grey; reconstruction, pink).
@@ 4.2  Results
GM
0.01581
FDM
0.00299
Arc-GM
0.01385
Arc-FDM
0.01345
Elastic-SRV-FDM
0.00424
Soft-SRV-FDM
0.00481
Arc-Elastic-SRV-FDM
0.00841
Arc-Soft-SRV-FDM
0.00794

### v2

@@ 1  Introduction
Functional data analysis (FDA) presents then a promising alternative by modeling curves as smooth functions within a functional space ramsay2005functional , treating the shape as a realisation of a stochastic process rather than a set of discrete coordinates. Srivastava and Klassen (2016)SrivastavaKlassen2016 motivate the need for rigorous statistical tools to analyse functions and shapes while handling invariances such as translation, rotation, scaling, and re-parameterisation. Their geometric framework unifies registration and shape analysis, enabling tools such as functional principal component analysis (FPCA) for vector-valued functions. Wu et al. (2024)wu2024shape presents a framework that integrates geometric shape analysis with functional data methods to better capture structural features of curves, such as peaks and valleys, beyond what traditional based approaches offer. The framework developed tools for shape fitting, shape-fPCA, and shape regression, extending these ideas to manifold-valued functions and demonstrating their effectiveness on real and simulated datasets.
@@
This work is organised as follows. Section 2 details the methods based on the GM and FDM frameworks. Section 3 presents a simulation study to compare FDM with its classical counterpart GM. Section 4 describes the application to 3D kangaroo landmark data. Section 5 provides the discussion for both the simulation study and empirical data. Section 6 concludes.
@@ 3.2  Simulation results
Figure 3 shows differences in classification performance, where functional data–driven approaches achieved higher predictive accuracy and better stability between folds. Figure 4 illustrates that functional pipelines reached the 95% variance threshold with fewer components than coordinate-based methods (Table 1). The corresponding reconstruction plot (Figure 6) confirmed that the reconstructed shapes using the retained components preserved the curved structural information, thus validating the use of the 95% cutoff. Both elastic and soft-aligned pipelines outperformed the purely coordinate-based and standard functional approaches. The elastic variants generally achieved good accuracy and efficient variance capture, but their performance showed greater variability across folds, indicating reduced stability. In contrast, soft-aligned pipelines offered consistently better accuracy with reduced variability and clearer group separation in PC space (Figure 7).
@@
GM 60
FDM 14
Arc-GM 41
Arc-FDM 15
Elastic-SRV-FDM 58
Soft-SRV-FDM 13
Arc-Elastic-SRV-FDM 51
Arc-Soft-SRV-FDM 13
@@
GM 0.01994 (0.00034)
FDM 0.00016 (0.00081)
Arc-GM 0.02089 (0.00363)
Arc-FDM 0.00425 (0.00439)
Elastic-SRV-FDM 0.00261 (0.00112)
Soft-SRV-FDM 0.00204 (0.00081)
Arc-Elastic-SRV-FDM 0.00197 (0.00126)
Arc-Soft-SRV-FDM 0.00192 (0.00120)
@@
Figure 6: Reconstruction landmark configuration of specimen 1 via arc-elastic-SRV-FDM (original, grey; reconstruction, pink).
@@ 4.2  Results
GM 0.01581 (0.00001)
FDM 0.00299 (0.00010)
Arc-GM 0.01385 (0.00068)
Arc-FDM 0.01345 (0.00069)
Elastic-SRV-FDM 0.00424 (0.00033)
Soft-SRV-FDM 0.00481 (0.00018)
Arc-Elastic-SRV-FDM 0.00841 (0.00056)
Arc-Soft-SRV-FDM 0.00794 (0.00042)
@@ 6  Conclusion
Disclosure Statement
During the preparation of this manuscript, the authors used ChatGPT (OpenAI) to improve the academic tone and enhance clarity of presentation. The tool was also used to reduce potential coding errors, improve readability, and ensure tidy code. All methodological design, analyses, and scientific interpretations were conducted solely by the authors. All content generated or suggested by the tool was critically reviewed, edited, and validated by the authors, who take full responsibility for the final version of the manuscript.